Q2 if Z is a standard normal varianle find the probability
P(-0.73 < Z < 2.27)
a. -0.2211
b. 0.4884
c. 0.7557
d. 0.2211
e. 1.54

Q3 Assume that women ties are normally distributed with a mean of 63.6 inches in a standard deviation of 2.5 inches if 60 women are randomly selected find the probability that they have a mean height between 62.9 inches and 64.0 inches
-assume that women's heights are normally distributed with a mean of 63.6 inches and a standard deviation of 2.5 inches if 60 women are randomly selected find the probability that they have a mean height between 62.9 inches and 64.0 inches
a. 0.9426
b. 0.1738
c. 0.8773
d. 0.0574
e. 0.1228
Answer #1

2)

P(-0.73<Z<2.27)=0.9884-0.2327=0.7557

3)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 63.6
std deviation   =σ= 2.500
sample size       =n= 60
std error=σ=σ/√n= 0.32275
probability =P(62.9<X<64)=P((62.9-63.6)/0.323)<Z<(64-63.6)/0.323)=P(-2.17<Z<1.24)=0.8924-0.015=0.8773